one mole of nitrogen and 3mole pcl5 are placed in a 100litre vesselheated to 227degreecelsius.The equlibrium pressure is 2.05 atm. assuming ideal behaviour,calculate kpfor the reaction?
can i get the solution please.

Grade:11

1 Answers

Vikas TU
7 years ago
Pcl5 would dissociate to Pcl3 + cl2 and P (N2 ) would not change it would remain same.
Therfore,
       Pcl5  ---------------->  pcl3   + cl2
       3 mol                     0             0
     3-x                           x          x
at eqm. total pressure is given as: 2.05
apply PV = nRT for total contents.
2.05*100 = n*0.0821*500
calculate from here n,
and then,
n = 3-x + x + x + 1
take x from here,
Then write easily Kc,
and for Kp = mole fracn.*Kc^1